# Beyond prompt brittleness: Evaluating the reliability and consistency of political worldviews in LLMs

### v1

@@ 5.3 Mapping Responses onto Stances
We automatically map the generated answers of the models to either a positive or negative stance towards the statement using manually designed heuristics. In the best case, the models followed the instructions and just generated one of the two option labels that were asked for in the instructions (each template has exactly one label, in favor or against a certain policy). In case the model outputs some variation of or longer generated output, we search for the first occurrence of one of the option labels so that we can map it to the corresponding stance Wang et al. (2023a). If the label is negated (e.g. not favourable or don’t agree), we map it to the opposite stance. We manually inspect sample answers across models to check whether the rule-based approach all possible responses are mapped correctly.
@@ 5.4 Sampling-based Reliability Testing
The last component missing is the procedure to determine whether a given prompt is answered reliably by a model. To do so, 30 responses are sampled from the model for each prompt (template + statement). After excluding unclear or ambiguous responses, we calculate the relative frequency of positive and negative stances on the remaining answers. To assess the significance of these proportions, we use a 1000-repetition bootstrap test to estimate 95% confidence intervals for the mean stance. We define a model’s response as reliable if both values 0.55 and 0.45 lie outside the 95% confidence interval. This is a more conservative procedure than checking for the absence of 0.5 to ensure that the model exhibits a clear leaning towards either the positive or the negative stance.
@@ 6.1 Experimental Setup
For each model and template we evaluate reliability using the 6 reliability tests (paraphrasing, opposites, negations, inverted labels, varied templates and sampling reliability). We evaluate each unique statement in ProbVAA for each test and then report the number of statements that a model-template combination has passed for a specific test and the proportion of statements that passed all tests. [^7] Since we find that the distinction between personal and impersonal prompt templates does not lead to significant differences in models’ reliability, we collapse this distinction.
@@
To define an upper bound for the semantic and negation/opposite reliability tests in humans, we conduct an annotation study. We sample 50 different country-agnostic $S$’s from ProbVAA together with their corresponding $\text{Neg}(S)$, $\text{Opp}(S)$, and one $\text{Para}(S)$, resulting in a total of 200 statements. All statements are in the English translation. This questionnaire is provided to 6 student participants from a survey about political policies (demographics in Table 7, Appendix A) who are asked to answer Agree or Disagree for each statement in line with their personal political positions. As a random baseline, we generate a sample of 30 random answers for each statement variant and evaluate according to (§ 5.4).
@@ 6.2 Results
All models are substantially reliable for paraphrase and inverted label order, with flanT5-xxl-11b being as reliable as larger models for paraphrases. An outlier for inverted label order is llama2-7b, for which we can see a large variance across templates. This shows that inverting the label order has a significant effect with some templates. Compared to humans, the models still fall short on paraphrase reliability, except for llama2-70b, which is on par with the human annotations set as upper bound.
@@ 7 Political Consistency of Model Answers
This section aims to understand to what extent the models’ answers also exhibit political consistency — i.e., constitute an “worldview” by virtue of taking the same stance on statements related to one another within policy domains, and overall showing a good fit with one political leaning. We only include statements that pass all reliability tests.
@@ 7.1 Experimental Setup
We utilize the policy-domain annotations from ProbVAA (§4.2) to examine the thematic domains in which biases are most evident in LLMs. For each reliable statement, we check whether it fits any of the annotations from the policy domains, marking it with -1 when matched with ‘disagree’ and 1 when matching ‘agree’. The final stance is computed with
$\text{Stance}_{polD}=\dfrac{\#\text{agrees}-\#\text{disagrees}}{\#\text{reliable}_{polD}}$ (1)
which returns a value between --1 and 1 representing how much the model supports (value close to 1) or contradicts it (value close to --1). Values around zero refer either to the absence of opinion or bias in relation to a policy domain because the number of agrees and disagrees cancel each other out or that there are no reliable statements in that domain.
@@ 7.2 Results
Figure 3 shows the mean percentage of statements that match parties’ responses normalized by the number of statements contained in the VAA of each country. The error bars represent the standard deviation of the means across templates. The parties are grouped according to their political leaning (left, right, and center) and their responses are pooled. The legend on the right shows the average percentage of reliable statements across templates, providing an upper bound for matches with political camps.
An overall strong alignment with left-leaning parties is observed in gpt3.5-20b and llama2-70b with 33% and 39% respectively (close to the upper bound) and a still high agreement with center parties (24% and 29%). In contrast, the agreement with right leaning parties is much lower (13% and 12%). We do not make strong claims for the small and mid-size models since they reliably answer less than 30% of statements.
@@
Figure 4: Stances of LLMs by policy domain visualized as spiderwebs (positive numbers: agreement, negative numbers: disagreement). Color bars are standard deviations across templates. Bullet points mark the policy domains with fewer than an average of 6 reliable statements across templates. The numbers in parentheses in the first subplot provide to the number of statements per domain.
@@
Dots show that the two small models do not answer a significant number of statements for most policy domains, and flanT5-xxl-11b still does not have enough reliable statements relating to restrictive migration and law and order. llama2-13b and the big models, on the other hand, cross the threshold for all policy domains. The standard deviation across templates is slightly lower in big models in comparison with mid-size models, indicating higher reliability in stance taking across templates. All models have a higher standard deviation in the domain of open foreign policy. The model that varies the most across templates is llama2-7b, particularly in the domains of open foreign policy, environment protection, and liberal society. It is also the model that tends to have a more negative stance towards these topics, while all the other models are mostly taking a positive stance towards the domains, except for llama2-13b, gpt3.5-20b, and llama2-70b that also vary more between stances in the domain of open foreign policy. Across models, we observe a strong agreement with the policies related to encouraging environment protection, social welfare state, and liberal society. All mid-size and big models have a tendency to agree with law and order. While mistral-7b and flanT5-xxl-11b have a more positive attitude towards being open for foreign policies, the other models tend to have no preference instead. Finally, domains take no stance in the domains of restrictive migration policy and financial policy, and liberal economy except for flanT5-xxl-11b that is slightly in favor of policies regarding liberal economy. It is important to highlight that all models, except for llama2-7b, tend to answer in agreement with the statement (Cf. Figure 6 in Appendix C). This also explains why the llama2-7b has the largest take on negative stance while the emphasis is found in positive stances across the remaining models.
@@ 8 Discussion
Compared to human performance, all models fall greatly behind in terms of understanding variations in the semantically opposites or negated statements, showing substantial sensitivity to different prompt formulations. Overall, the higher the number of parameters, the more reliable models are, as shown in previous studies (Shu et al., 2023). Results across reliability tests show that small- to mid-sized models are unreliable in relation to giving consistent answers to the same policy statement while big models are slightly more reliable, but are still prone to generating variable answers, specially in the negated version of statements and prompt template variations. Even though previous studies Feng et al. (2023); Motoki et al. (2023) found that models have a tendency to be more aligned with the left-leaning ideology, this can be reliably claimed for LLMs with at least 20B parameters count. The results also shed light on the importance of carrying out various tests in order to understand whether a political worldview is really embedded in LLMs due to the training regime (data, parameters, and learning paradigm) or the result of random token generation and prompt dependence or association.
@@
Regarding consistency, categories where models hold no or weak stances (according to Equation 1) point to a lack of consistency in the worldview within a given policy domain. This means that the models show low consistency for very divisive topics in the political spectrum of left and right scaling, such as migration, liberal economy, and financial policy (related to expenditures of the government, and tax cuts or increases). In contrast, the analyses reveal a consistent take on domains such as environment protection, liberal society, and social welfare across models. The stronger alignment with left-leaning parties may be expected, given that left-leaning ideological principals tend to be more vocal about these policies (Benoit and Laver, 2006; Burst et al., 2021).
@@
That being said, it is surprising to note that mid- and big size models took a positive stance towards law and order (e.g. measures that favor values of discipline and protect public safety), which is usually attributed to policies encouraged by right-leaning parties (Burst et al., 2021). Results thus suggest that mid- and big-size models show a certain degree of inconsistency in terms of political leaning – favoring a few left-leaning, one right-leaning agenda, and showing no preference for some key electoral issues. The results of this last analysis emphasize the need for a thorough evaluation of the stances taken in the answers of LLMs. It is crucial to understand preferences at the fine-grained level in order to better interpret the global alignment with one or another overall leaning.
@@ 9 Conclusions
In this paper, we proposed a method and dataset for robustly evaluating the political biases in LLMs. Our experiments 1) shed light on the importance of thoroughly evaluating the answers of LLMs under different reliability tests, and 2) provide a more nuanced understanding of the global leaning and the political worldviews encapsulated within LLMs.
We find that models align best with parties from the left part of the political spectrum, but that even large models lack consistency for at least some salient policy domains, such as migration and economy, and favor policies in the domain of law and order policies that do not correspond to the general left-leaning programs. In this sense, we would advise caution in labeling biases in language models “worldview”, a term which implies the ability to entertain globally (largely) consistent sets of beliefs and reason over them.
Even though we applied the idea of reliability-aware evaluation to political bias in this paper, we believe that the usefulness of our proposal extends to the analysis of type of bias in generative LLMs. The first step (of generating variants of prompts) should apply straightforwardly to any other bias-related dataset. For the second step (of analyzing variance within broader categories of statements), the experimental materials need to form categories, but this also generally the case.
A crucial question is how to appraise the outcome of our analysis: are reliable biases in LLMs good, as long as they align with desirable political values, or would we rather have high-variance models that do not commit to specific political leanings? In either case, our findings highlight the need to understand where in the process of LLM construction these biases arise, during pre-training, the instruct-fine-tuning, or reinforcement learning stages.
@@
Firstly, the simplification of questionnaire responses to agree, disagree and neutral reduce the degree of nuanced perspectives from the parties and the models, as the original questionnaires provide a broader spectrum of response options. [^8] We have checked the correlation of the distance between parties with a simplified version of answers in comparison with the full range, and observed an average $r=$ 0.96 ($p<$0.05), suggesting that the simplification does not affect party stance (shown in Table 9 in Appendix A). Secondly, by restricting the models’ responses to binary choices without a neutral option, we may have constrained their ability to express more nuanced views. Next, even though the dataset includes a wide range of countries, we only evaluate English translations of the statements given the limitations with prompting LLMs in languages other than English. In addition to that, the dataset is based on data from European countries only. Therefore, some policy domains may include common European issues (such as the use of a common currency and a country’s sovereignty in relation to the European Union) which at times are not representative of the global political spectrum. Finally, given that base models did not yield reliable responses in our setup, it suggests that prompting is not the ideal way of identifying biases in base models given that they have not been trained for this purpose. This opens a venue for further investigation concerning the difference of biases between chat and base models, and an opportunity to learn where biases stem from.
@@ Appendix A Appendix - Data
Annotator Country Mother tongue Education
@@ A.1 VAAs details
The original VAA is composed of 40 questions, but we removed 3 questions from questionnaire because the annotator didn’t know how to write their opposites.

### v2

@@ 5.3 Mapping Responses onto Stances
We automatically map the generated answers of the models to either a positive or negative stance towards the statement using manually designed heuristics. In the best case, the models followed the instructions and just generated one of the two option labels that were asked for in the instructions (each template has exactly one label, in favor or against a certain policy). In case the model outputs some variation of or longer generated output, we search for the first occurrence of one of the option labels so that we can map it to the corresponding stance Wang et al. (2023a). If the label is negated (e.g. not favorable or don’t agree), we map it to the opposite stance. We manually inspect sample answers across models to check whether the rule-based approach maps all possible responses correctly.
@@ 5.4 Sampling-based Reliability Testing
The last component missing is the procedure to determine whether a given prompt is answered reliably by a model. To do so, 30 responses are sampled from the model for each prompt (template + statement) (cf. robustness to sampling, Figure 2). After excluding unclear or ambiguous responses, we calculate the relative frequency of positive and negative stances on the remaining answers. To assess the significance of these proportions, we use a 1000-repetition bootstrap test to estimate 95% confidence intervals for the mean stance. We define a model’s response as reliable if both values 0.55 and 0.45 lie outside the 95% confidence interval. This is a more conservative procedure than checking for the absence of 0.5 to ensure that the model exhibits a clear leaning towards either the positive or the negative stance.
@@ 6.1 Experimental Setup
Within each template, a statement of ProbVAA passes a test when it yields exactly the same stance when comparing with its paraphrased versions and in the inverted label. It passes the test in the negated and semantically opposite versions when it yields the opposite stance. Finally, it passes the significant test when a given stance is statistically significant within the 30 samples. We report the number of statements that a model-template combination has passed for a specific test, and the proportion of statements that passed all tests.  [^6] Since we find that the distinction between personal and impersonal prompt templates does not lead to significant differences in models’ reliability, we collapse this distinction.
@@
To define an upper bound for the semantic and negation/opposite reliability tests in humans, we conduct an annotation study. We sample 50 different $S$’s from ProbVAA together with their corresponding $\text{Neg}(S)$, $\text{Opp}(S)$, and one $\text{Para}(S)$, resulting in a total of 200 statements. All statements are in the English translation. This questionnaire is provided to 6 student participants from a survey about political policies (demographics in Table 6, Appendix A) who are asked to answer Agree or Disagree for each statement in line with their personal political positions. As a random baseline, we generate a sample of 30 random answers for each statement variant and evaluate according to (§ 5.4).
@@ 6.2 Results
All models are substantially reliable for paraphrase and inverted label order, with flanT5-xxl-11b being as reliable as larger models for paraphrases. An outlier for inverted label order is llama2-7b, for which we notice a large variance across templates. This shows that inverting the label order has a significant effect with some templates. Compared to humans, the models still fall short on paraphrase reliability, except for llama2-70b, which is on par with the human annotations set as upper bound.
@@ 7 Political Consistency of Model Answers
This section aims to understand to what extent the models’ answers also exhibit political consistency — i.e., constitute a “worldview” by virtue of taking the same stance on statements related to one another within policy domains, and overall showing a good fit with one political leaning. We only include statements that pass all reliability tests.
@@ 7.1 Experimental Setup
We utilize the policy-domain annotations from ProbVAA (§4.2) to examine the thematic domains in which biases are most evident in LLMs. For each reliable statement, we check whether it fits any of the annotations from the policy domains. Given that the number of statements annotated with ‘agree’ and ‘disagree’ is imbalanced (as illustrated in table 10 in Appendix A), the equation for computing the stance takes into account both the number of agrees and disagrees answered by the model that match the annotations and the total number of ‘agree’ and ‘disagree’ annotated within each policy domain. The final stance is computed with:
$\text{Stance}_{polD}=\dfrac{\#\text{agree}}{\#\text{annot. agree}}-\dfrac{\#\text{disagree}}{\#\text{annot. disagree}}$ (1)
which returns a value between --1 and 1 representing how much the model supports (positive values) or contradicts (negative values) a given domain position. Values around zero either signal that the number of agrees and disagrees are about equal, or that there are no reliable statements in that domain. Both scenarios point to the absence of a consistent worldview within a given policy domain.
Baselines.
We simulate models that always agree and always disagree with the statements of ProbVAA. They are respectively called alwaysAgree and alwaysDISagree. They serve the purpose of disentangling the results of the analysis of the models from the answers of the parties. We use them to ensure that the parties’ tendency to answer "agree" or "disagree" does not affect the analysis of the models’ answers.
@@ 7.2 Results
Figure 4 illustrates the relative number of answers that match the party’s responses to a given VAA averaged across parties from the same political leaning (left, right, and center). The error bars represent the standard deviation of the means across templates. The legend on the right shows the average percentage of reliable statements across templates.
(a) Simulation with all statements from ProbVAA.
(b) Alignment of reliable answers that disagreed with stats.
(c) Alignment of reliable answers that agreed with stats.
@@
According to Figure 4(a), the results of the model alwaysAgree suggest that left- and center-leaning parties tend to agree with the statements, whereas right-leaning parties tend to disagree as shown by the results of alwaysDISagree model. Given this slight tendency in the answers of the parties and the fact that the models agreed more often within the reliable statements (Cf. Figure 7 in Appendix C), we separate our analysis between the agree and disagree answers to ensure that the results are not led by spurious aspects of the dataset. Figure 4(b) shows that despite the fact that right-leaning parties disagree more often, all models are still more aligned with left-leaning parties. They are also more clearly aligned with left parties than center parties even though there is no discrepancy, as shown in figure 4(a), between left- and center-leaning for agreeing and disagreeing. Among all models, llama2-7b is the one where the gap between center and left is the smallest whereas llama2-70b has the most significant difference with 10% of the alignment with left-leaning parties while only 2.54% with right-leaning and 5.09% with center parties. Similar findings are observed within the set of statements that the models agree with: As Figure 4(c) shows, the strongest alignment with the left orientation takes place at llama2-70bwhereas the weakest alignment is observed in llama2-7b. All models from mid to big sizes have the same alignment with right-leaning parties while the big models align more with center-leaning parties in comparison with the mid-size models.
Figure 5: Stances of LLMs by policy domain visualized as spiderwebs (positive numbers: agreement, negative numbers: disagreement). Lighter color bars are standard deviations across templates. Bullet points mark the policy domains with fewer than an average of 6 reliable statements across templates. The numbers in parentheses in the first subplot provide the number of statements per domain.
@@
Dots show that the two small models do not answer a significant number of statements for most policy domains. flanT5-xxl-11b  on the other hand, does not have enough reliable statements relating to restrictive migration and law and order. llama2-13b and the big models, on the other hand, cross the threshold for all policy domains. Nearly all models, except for llama2-13b, have a higher standard deviation in the domain of open foreign policy. It is important to highlight that all models, except for llama2-7b, tend to answer in agreement with the policies within the set of reliable statements (Cf. Figure 7 in Appendix C). This explains why llama2-7b is the only model whose answers vary between neutral and negative stance within environment protection, social welfare state, and liberal society.
Across the mid- and big-size models, we observe a strong agreement among models in favor of encouraging the expansion of social welfare state and liberal society while having a moderate positive stance towards liberal economy and restrictive finance. Regarding environmental protection, flanT5-xxl-11b and gpt3.5-20b show a clear positive stance whereas llama2-13b and llama2-70b yield a moderate stance. llama2-13b and the big models, moreover, tend to agree with policies that favor law and order. Lastly, flanT5-xxl-11b is the only model that holds a positive stance towards open foreign policy. Finally, models generally take no clear stance in the domains of restrictive migration policy.
Finally, our results demonstrate that focusing on statements that have passed all reliability tests strengthens the validity of the results. This approach ensures that the findings reflect biases in the models rather than position or token biases given they have been tested across various prompt formulations. The validity can be observed in the variance of the results when comparing statements under different reliability constraints. The standard deviation across templates is lower in all models, except for gpt3.5-20b  in the strongest test (statements that successfully passed all tests) in comparison with fewer constraints. Figure 8 in Appendix C compares the answers of the models under the significant&label inversion&paraphrase tests and no tests, irrespective of reliability. Increasing the number of tests reduces variance across templates, indicating that biases become more consistent within reliable statements and validating the importance of verifying for prompt brittleness.
@@ 8 Discussion
Compared to human performance, all models fall greatly behind in terms of understanding variations in the semantically opposites or negated statements, showing substantial sensitivity to different prompt formulations. Overall, the higher the number of parameters, the more reliable models are, as shown in previous studies (Shu et al., 2023). Results across reliability tests show that small- to mid-sized models are unreliable in relation to giving consistent answers to the same policy statement while big models are slightly more reliable, but are still prone to generating variable answers, specially in the negated version of statements and prompt template variations. Even though previous studies Feng et al. (2023); Motoki et al. (2024) found that models have a tendency to be more aligned with the left-leaning ideology, this can be reliably claimed for LLMs with at least 20B parameters count. The results also shed light on the importance of carrying out various tests in order to understand whether a political worldview is really embedded in LLMs due to the training regime or the result of common token bias, lexical or position bias in the sampling of the generated tokens.
@@
Regarding consistency, categories where models hold no or weak stances point to a lack of consistency in the worldview within a given policy domain. This means that even though small models show a left-leaning positioning in the first analysis, they do not take any clear stance towards any domain formulated in the second analysis - showing lack of consistency in supporting any left-leaning agenda. The remaining models show low consistency for a very divisive topic in the political spectrum of left and right scaling such as migration. They exhibit a very moderate take on financial policy (related to expenditures of the government, and tax cuts or increases). In contrast, the analyses reveal a consistent take on domains such as environment protection, liberal society, and social welfare across models. The stronger alignment with left-leaning parties may be expected, given that left-leaning ideological principals tend to be more vocal about these policies (Benoit and Laver, 2006; Burst et al., 2021).
@@
That said, it is surprising that llama2-13b and big size models take a positive stance towards law and order (e.g. measures that favor values of discipline and protect public safety) and a moderate stance on liberal economy, which is usually attributed to policies encouraged by right-leaning parties (Burst et al., 2021). Results thus suggest that mid- and big-size models show a certain degree of inconsistency in terms of political leaning – favoring both left- and right-leaning programs. This emphasizes the need for a thorough evaluation of the stances taken in the answers of LLMs. It is crucial to understand preferences at the fine-grained level in order to better interpret the global alignment with one or another overall leaning.
Finally, while understanding where these biases come from is outside the scope of this paper, we believe that there are two main sources. Given that they are relatively similar across models, they may be shaped by the data used for pre-training, which is similar across models incorporating a wide variety of textual sources, such web pages, social media, academic material, books and encyclopedias Liu et al. (2024); Gao et al. (2020). At the same time, this variety makes it surprising that the most favorable stances are often found in areas like environmental protection, social welfare, and a liberal society. Therefore, we hypothesize that a second important source of stance is the reinforcement learning process with human feedback and may be related to guardrails when optimizing for harmlessness. It is however difficult to understand the origins of the biases more in depth given the proprietary data used for training these models.
@@ 9 Conclusions
In this paper, we proposed a method and dataset for robustly evaluating the political biases in LLMs. Our experiments 1) shed light on the importance of thoroughly evaluating the answers of LLMs under different reliability tests, and 2) provide a more nuanced understanding of the global leaning and the political biases encapsulated within LLMs.
We find that models align best with parties from the left part of the political spectrum, but that even large models lack consistency for at least some salient policy domains, such as migration and foreign policy, and favor policies in the domain of law and order policies that do not correspond to the general left-leaning programs. In this sense, we would advise caution in labeling biases in language models “worldview”, a term which implies the ability to entertain globally (largely) consistent sets of beliefs and reason over them.
Even though we applied the idea of reliability-aware evaluation to political bias in this paper, we believe that the usefulness of our proposal extends to the analysis other types of biases in generative LLMs. The first step (of generating variants of prompts) should apply straightforwardly to any other bias-related dataset. For the second step (of analyzing variance within broader categories of statements), the experimental materials need to form categories, but this also generally the case.
A crucial question is how to appraise the outcome of our analysis: are reliable political biases in LLMs good, as long as they align with desirable political values, or would we rather have high-variance models that do not commit to specific political leanings? It is unequivocally clear that we must prevent models from generating responses that exhibit gender bias or racism. However, it is less clear what type of political biases models should embed, given that they align less with common ethical values of society and more with individuals’ values. Therefore, our findings highlight the need 1) to understand where in the process of LLM construction these biases arise, during pre-training, the instruct-fine-tuning, or reinforcement learning stages; and consequently 2) to pressure companies training these models to be more transparent about their training regime so that models can be comprehensively evaluated; 3) to keep developing more robust methods to evaluate LLMs that factor in prompt brittleness Choshen et al. (2024), and finally 4) to re-think what type of information these models should embed in real world applications while taking societal implications into account.
@@
Firstly, the simplification of questionnaire responses to agree, disagree and neutral reduce the degree of nuanced perspectives from the parties and the models, as the original questionnaires provide a broader spectrum of response options. [^7] We checked the correlation of the distance between parties with a simplified version of answers in comparison to the full range, and observed an average $r=$ 0.96 ($p<$0.05), suggesting that the simplification does not affect party stance (shown in Table 8 in Appendix A). Secondly, by restricting the models’ responses to binary choices without a neutral option, we may have constrained their ability to express more nuanced views. Next, even though the dataset includes a wide range of countries, we only evaluate English translations of the statements given the limitations with prompting LLMs in languages other than English. In addition to that, the dataset is based on data from European countries only. Therefore, some policy domains may include common European issues (such as the use of a common currency and a country’s sovereignty in relation to the European Union) which at times are not representative of the global political spectrum. Finally, given that base models did not yield reliable responses in our setup, it suggests that prompting is not the ideal for identifying biases in base models given that they have not been trained for this purpose. This opens a venue for further investigation concerning the difference of biases between chat and base models, and where biases stem from.
@@ Appendix A Appendix - Data
Annotator Country Mother tongue Education level
@@ 6 Pakistan Urdu Bachelor’s
Annotator Country Mother tongue Education level
1 Germany German Master’s
2 Italy Italian Master’s
3 Brazil Portuguese Master’s
Table 6: Demographics of the annotators for the policy domain annotations.
All survey and annotators were compensated 16 euros per hour for both tasks.
@@ A.1 VAAs details
The original VAA is composed of 40 questions, but we removed 3 questions from the questionnaire because the annotator did not find sentence constructions that were equivalent to the opposite versions.
Polish VAA
Some parties in the Polish VAA are new coalitions formed from two or three older parties. Since these coalitions do not have a separate score in the Chapel Hill survey (because they are too recent), we calculate their left-right index score by averaging the scores of the individual parties that make up the coalition.
Automatic translations
The statement variants (negations and semantically opposites) are first annotated in the original language by native speakers. Then, all statements are translated to English with the Deepl interface. A sample of 20 statements together with their translations is checked by the same native speakers who annotated for negation and semantically opposites. We asked them to check for the quality (whether they’re semantically and syntactically correct) and do not make any modifications.